How do you solve:
Determine the x-intercept and y-intercept of 3x-8y=12
Thank you

Answer by Alan3354(69443)

Determine the x-intercept and y-intercept of 3x-8y=12
set x = 0
y = -3/2 (y-intercept)
---------
at y = 0, x = 4 (x-intercept)